Give the following probabilities:、

P(success)=P(s)=0.7

P(favorable forecast s)= P (f/s)=0.85

P(favorable forecast /not pass)= P (f/n)=0.1

compute p(s/f) (s/u)and P(F)
Answer #1

Ans:

P(F)=P(f/s)*P(s)+P(f/u)*P(u)

=0.85*0.7+0.1*(1-0.7)=0.625

P(s/f)=P(f/s)*P(s)/P(F)

=0.85*0.7/0.625=0.952

P(u/f)=P(f/u)*P(u)/P(f)

=0.1*(1-0.7)/0.625=0.048